OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (4, 4) and (2, 6) (2, 4) and (4, 2) Answer The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is smaller than the slope of Set #2.

OpenStudy (anonymous):

Slope of set 1 = -1 slope of set 2 = -1 So they are same

OpenStudy (anonymous):

How about this one? Select the equation that corresponds with the given graph. This is a graph of a coordinate plane with a line. Two points of the line are zero comma five and negative five comma four. Answer y = -5x + 5 y equals negative 1 over 5 times x + 5 y equals 1 over 5 times x + 5 y = 5x + 5

OpenStudy (anonymous):

but (-5,4) doesnt lie on y = 5x +5 the eqn i got is 5y=x+25

OpenStudy (anonymous):

oh. u did write it properly in words.

OpenStudy (anonymous):

Yeeah, I did....